OpenStudy (anonymous):

Need help Finding the point on the terminal side of theta= -3pi/4 that has an x coordinate of -1.

OpenStudy (debbieg):

There's more than one approach to this, but if you know your unit circle values then I guess the simplest is this: \(\large \tan(3\pi/4)=y/x\) \(\large \tan(3\pi/4)=y/(-1)\) So set the value of \(\large \tan(3\pi/4)=y/(-1)\) and solve for y. Then you will have the two coordinates of the point. Alternatively, you can "draw the picture", and solve the triangle (you need to know what the angle a is). That should give you the y coordinate: 

OpenStudy (debbieg):